The total revenue function for two goods is given by the equation TR = 36 x-3x²+ 56y- 4y².find the number of unit of each goods which must be sold. if profit is to be maximize when the firm subject to a budget constraint 5 x + 10 y = 80in Lagrange multiplayer method
1
Expert's answer
2021-10-27T13:55:37-0400

Total revenue function:

"TR=36x-3x^2+56y-4y^2"

Budget constraint:

"5x+10y=80"

"x+2y=16"

"x=16-2y"

Substituting the value of x in the total revenue equation gives:

"TR=36(16-2y)+3(16-2y)^2+56y-4y^2."

Differentiating the above equation w.r.t y:

"\frac{\delta TR}{\delta y}=-72y+ 6(16-2y)+56+8y"

"=-104y+248=0"

"\implies" "y=2.39"

"x=16-2y"

"=16-2(2.385)=11.23"

"x=11.23"

Thus, the quantities of goods x and y to be sold are 11.23 and 2.39 respectively.